OpenStudy (anonymous):

How many solutions can be found for the equation 4z + 2(z − 4) = 3z + 11? None One Two Infinitely many

OpenStudy (anonymous):

A)None B) One C)Two D)Infinitely many

OpenStudy (anonymous):

please help?

OpenStudy (kidrah69):

Ok we do the same thing as the last problem

OpenStudy (anonymous):

ok

OpenStudy (kidrah69):

\[4z + 2(z − 4) = 3z + 11 \] we need to distribute first so... \[2(z − 4) \] you get \[4z+2z-8=3z+11\] combine like terms so what is... \[4z+2z=?\]

OpenStudy (anonymous):

6

OpenStudy (kidrah69):

Good so now we have... \[6z-8=3z+11\] combine like terms add 8 to both sides and get... \[6z=3z+19\] now combine the Zs \[6z+3z=?\]

OpenStudy (anonymous):

9

OpenStudy (kidrah69):

good what is \[19 \div 9=?\]

OpenStudy (anonymous):

2.1111

OpenStudy (kidrah69):

so the answer would be what??

OpenStudy (anonymous):

A?

OpenStudy (kidrah69):

yep :)

OpenStudy (anonymous):

thanks